OpenStudy (anonymous):

Given the equation y - 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. Answer -3/4 x + y = 10 3x - 4y = -40 y = 3/4x + 12 y = 3/4 x + 10

OpenStudy (callisto):

y - 4 =3/4 (x + 8) y =3/4 (x + 8) +4 = 3/4 (x) + (3/4)(8) +4 = 3/4 (x) + 6+4 =?